# Eigenstate Thermalization Hypothesis (ETH) for off-diagonal matrix elements in integrable spin chains

### v1

@@ I Introduction
A fundamental problem in the study of out-of-equilibrium quantum many-body physics is to understand how the unitary evolution of isolated systems leads to local equilibration and thermalization. Our current understanding of thermalization is based on the celebrated Eigenstate Thermalization Hypothesis [1, 2, 3, 4] (ETH). Let $H$ be a generic, i.e., chaotic, many-body Hamiltonian and let |$\ket{E_{i}}$ be its energy eigenstates $H\ket{E_{i}}=E_{i}\ket{E_{i}}$. According to ETH, the matrix elements of a generic local operator ${\mathcal{O}}$ between energy eigenstates take the form [1, 2, 3, 4]
$\bra{E_{i}}{\mathcal{O}}\ket{E_{j}}=O(\overline{E})\,\delta_{ij}+f_{\mathcal{O}}(\omega,\overline{E})\,e^{-S(\overline{E})/2}R_{ij}\,,$ (1)
@@
where $\overline{E}=\frac{1}{2}(E_{i}+E_{j})$ is the average energy of the eigenstates, $\omega=\left|E_{i}-E_{j}\right|$, $O(\overline{E})$ is a smooth function of the energy $\overline{E}$. In (1), $S(\overline{E})$ is the thermodynamic entropy at $\bar{E}$, $f_{\mathcal{O}}(\omega,\overline{E})$ is a smooth function of its arguments, and $R_{ij}$ is a random variable with mean $\overline{R_{ij}}=0$ and variance $\overline{R_{ij}^{2}}=1$. Eq. (1) implies that in the thermodynamic limit, since the entropy $S$ is extensive in the system size $L$, diagonal matrix elements converge to $O(\overline{E})$, i.e., they become smooth functions of the energy only. Moreover, the variance of diagonal matrix elements decays exponentially with $L$. Analogously, off-diagonal matrix elements are exponentially suppressed with increasing $L$. An intriguing recent result is that the statistics of the off-diagonal matrix elements can be understood in terms of Free Probability Theory [5, 6]. Crucially, Eq. (1) implies local thermalization after quantum quenches in generic, i.e., nonintegrable systems, at least for typical initial states [4]. The ETH scaling (1) has been checked numerically for both diagonal and off-diagonal matrix elements in nonintegrable systems by employing exact diagonalization methods [7, 8, 9, 10, 11, 12, 13, 14]. In integrable models, the presence of an extensive number of conserved charges leads to violations of (1), which have been investigated numerically [15, 16, 17, 18, 19, 20, 21, 22]. Refs. [23, 24] investigated the $ETH$ scaling of diagonal matrix element in interacting integrable systems by exploiting Algebraic Bethe Ansatz (ABA) results [23, 25] for correlation functions, which allow to reach much larger system sizes compared with exact diagonalization. Precisely, in Ref. [25] showed that the variance of diagonal matrix elements decays as $L^{-1}$, in contrast with (1).
@@
Very recently, Ref. [26] investigated off-diagonal matrix elements of both local and non-local operators in the Lieb-Liniger model, which describes a one-dimensional system of interacting bosons in the continuum. Ref. [26] showed that the finite-size scaling of off-diagonal matrix elements depends crucially on whether $|E_{i}\rangle$ and $|E_{j}\rangle$ (cf. (1)) are extracted from the same thermodynamic macrostate, i.e., statical ensemble. Precisely, if $|E_{i}\rangle,|E_{j}\rangle$ are in the same macrostate, off-diagonal matrix elements decay exponentially as [26]
$\bra{E_{i}}\mathcal{O}\ket{E_{j}}\propto\exp\!\left[-c^{\mathcal{O}}\,L\ln(L)-\widetilde{M}_{ij}^{\mathcal{O}}\,L\right],$ (2)
where $c^{\mathcal{O}}>0$ is a constant that depends both on the operator $\mathcal{O}$ and on the macrostate. Moreover, the probability distribution function of $\widetilde{M}_{ij}^{\mathcal{O}}$ is well described by a Fréchet distribution. Curiously, the leading term in the large $L$ limit in the exponent of (2) exhibits a logarithmic correction. Notice that the scaling (2) is similar to (1), except for the logarithmic correction and for the non Gaussian statistics of the $\widetilde{M}_{ij}^{\mathcal{O}}$. The matrix elements of local operators between eigenstates that are in different macrostates exhibit a faster decay in the thermodynamic limit as [26]
$\bra{E_{i}}\mathcal{O}\ket{E_{j}}\propto\exp\!\left(-d^{\mathcal{O}}\,L^{2}\right),$ (3)
@@
Quite generically, the PDFs are well described by Gumbel distributions, in contrast with Ref. [26]. This confirms the non Gaussian nature of the statistics of off-diagonal matrix element in integrable spin chains. Next, we consider two-spin operators. We focus on thermal macrostates constructed from eigenstates that do not contain strings. In contrast with one-spin operators, for which we can reach $L\sim 500$, for two-spin operators we only provide numerical data for $L\sim 60$. Despite that, our numerical data confirm the scenario observed for one-spin operators.
@@ II Heisenberg spin chain
The $XXX$ chain Eq. 5 is a paradigmatic example of Bethe Ansatz solvable many-body system [37, 36, 38, 39]. In the Bethe ansatz treatment of the $XXX$ chain the ferromagnetic state with all the spin up is a “vacuum” state, whereas overturned spins are treated as quasiparticle excitations. Each quasiparticle is labeled by a complex number $\lambda_{i}$, which is called rapidity. Any eigenstate in the sector with fixed number $M$ of quasiparticles, i.e., fixed magnetization, is identified by a set of $M$ rapidities $\lambda_{i}$. The rapidities satisfy nontrivial quantization conditions, encoded in a set of nonlinear coupled algebraic equations as
@@
The Bethe equations (6) admit $2^{L}$ sets solutions containing both real and complex rapdities. Complex rapidities correspond to bound states of quasiparticles. Although it is in general challenging to obtain the complex solutions of (6), it was already recognized by Bethe [37] that in the thermodynamic limit $L\to\infty$ some crucial simplifications occur. Precisely, the vast majority of complex rapidities form “simple” structures in the complex plane known as Bethe strings. Solutions of (6) in the same string have the same real part (string center) and imaginary parts that differ by $\eta=\mathrm{i}$. A string of length $n$ corresponds to a bound states of $n$ elementary excitations. In generic integrable systems, strings of arbitrary length are allowed.
Within the framework of the string hypothesis, let us denote the rapidities as $\lambda^{j}_{n,\gamma}$, with $\gamma$ labeling the different strings of size $n$, i.e., that have different string centers, and $j$ labeling the rapidities in the same string. For large $L$ one has for that
@@
where $\delta_{n,\gamma}^{j}$ are the so-called string deviations, which account for the fact that the string hypothesis holds only in the limit $L\to\infty$. In (8) $\lambda_{n,\gamma}$ is the string center, and it is real. Real solutions of the Bethe equations correspond to $n=1$. For the vast majority of the eigenstates of the $XXX$ chain string deviations are suppressed exponentially as $\delta_{j,\gamma}^{j}=\mathcal{O}(e^{-\alpha L})$ (see Ref. [40] for a detailed study of string deviations in the $XXX$ chain). In particular, string deviations are not expected to affect the thermodynamic behavior of the model. By employing (8) in Eq. 6, after taking the limit $\delta_{n,\gamma}^{j}\to 0$ and considering the logarithm of both members of the Bethe equations, one can write a set of equations for the string centers $\lambda_{n,\gamma}$. One obtains the so-called Bethe-Gaudin-Takahashi equations [36] as
@@
$\Theta_{n,m}=\theta_{|n-m|}+\!\!\!\!\!\!\sum_{r=1}^{(n+m-|n-m|-1)/2}\!\!\!\!\!\!\!\!\!\!\!\!\!2\theta_{|n-m|+2r}+\theta_{n+m}$ (10)
@@
$\Theta_{n,m}=\sum_{r=1}^{n-1}2\theta_{2r}+\theta_{2n}$ (11)
if $n=m$. In Eqs. 9, 10 and 11 we defined $\theta_{n}(x)=2\arctan(x/n)$. Let us denote by $s_{n}$ the number of strings of length $n$ in the eigenstate. Thus, one has the total number of particles as $M=\sum_{j=1}^{M}js_{j}$. In (9), $I_{n,\gamma}$ are the Bethe-Gaudin-Takahashi (BGT) quantum numbers, and are integer or half-integer numbers for $L-s_{n}$ odd or even, respectively. For the $XXX$ chain the BGT quantum numbers satisfy the bound as [36]
@@ II.1 Macrostates and observables
In the following section, in particular, we will study the statistics of the off-diagonal matrix elements of $E_{i}^{(11)}$ and $E_{i}^{(11)}E_{i+1}^{(22)}$, taken as prototypical examples of one and two-spin operators, respectively. The operators (16) are particularly convenient to compute in the $XXX$ model (5), by virtue of the exact results of Refs. [30, 31]. Indeed, a major breakthrough [30, 31] is that it is possible to rewrite the operators (16) in terms of the matrix elements of the so-called transfer matrix of the $XXX$ chain [38], which is at the heart of the ABA solution of the model. Then, by using the known action of the transfer matrix on the eigenstates of (5), it is possible to obtain the matrix elements. We review the main results of Refs. [30, 31] in Appendix A, reporting the explicit expression for the matrix elements of the operators in (16). Unlike exact diagonalization approaches, the computational cost to evaluate matrix elements via the ABA results reported in appendix A scales exponentially only in the size $\ell$ of the support of the operator (here we restrict ourselves to $\ell=1,2$), but polynomially as $O(M^{\ell})$ in the number of particles $M$. This allows us to reach chain sizes $L\sim 500$ which are prohibitively large for exact diagonalization. Still, we should mention that the numerical implementation of the ABA formulas for the matrix elements is nontrivial, because they are derived for an inhomogeneous version of the $XXX$ chain. The homogeneous limit that is required to recover (5) introduces fictitious singularities that have to be removed. Furthermore, the presence of string solutions, introduces additional singularities, which have to be dealt with by carefully taking the limit of vanishing string deviations. For the operators with support on more than one site, treating the homogeneous limit and the limit of vanishing string deviations at the same time is very challenging. For this reason, as in Ref. [24], we only consider off-diagonal matrix elements of $E_{i}^{(11)}E_{i+1}^{(22)}$ between eigenstates without bound states. For one-spin operators, since the homogeneous limit is trivial, the limit of vanishing string deviations can be performed analytically. Thus, for one-spin operators we are able to investigate the effect of bound states on the finite-size scaling of matrix elements.
@@ III Numerical results
Figure 1: Logarithm of the matrix elements for the one-spin operator $E^{(11)}_{j}$ with $L=56,\,M=14$ as a function of the energy densities $E_{i}/L$ of the two eigenstates (diagonal elements not reported).
Figure 2: Scaling of the logarithm $M_{ij}$ (cf. Eq. 17) of the matrix elements of the one-spin operator $E_{i}^{(11)}$ as a function of $E_{i}-E_{j}$, with $E_{i},E_{j}$ the energy of the two eigenstates. (a) Scatter plot of the logarithm $M_{ij}$ rescaled by $L$ as a function of $(E_{i}-E_{j})\sqrt{L}$ for $L=56,80,120,160,$ and $240$ and fixed $M=L/4$. (b) Two-dimensional histogram of the rescaled logarithm $M_{ij}/L$ as a function of $(E_{i}-E_{j})/\sqrt{L}$.
@@
$M_{ij}=\ln\left|\bra{E_{i}}\mathcal{O}\ket{E_{j}}\right|^{2}=\ln\left|\mathcal{O}_{ij}\right|^{2}.$ (17)
In Fig. 1 we overview the behavior of $M_{ij}$ for the one-spin operator $E^{(11)}_{i}$ (cf. (16)) in the $XXX$ chain with $L=56$, $M=L/4$. The data shown in the figure are for the infinite-temperature macrostate. Precisely, by employing the method of Ref. [42] we generate an ensemble of $\sim 10^{3}$ eigenstates of the $XXX$ chain distributed with the infinite-temperature Gibbs probability. As discussed earlier, we fix the magnetization $M/L=1/4$ and consider only the eigenstates that correspond to solutions of the Bethe equations without infinite rapidities. On the two axes we show the energy density $E_{1}/L$ and $E_{2}/L$ of the eigenstates. Here we have $E_{1}/L,E_{2}/L\in[-0.3,-0.1]$. Notice that in the thermodynamic limit $L,M\to\infty$, the fluctuations of the energy density are suppressed as $1/\sqrt{L}$, i.e., the energy density peaks at the thermal value at the density $M/L=1/4$. Again, this reflects that in the thermodynamic limit all the eigenstates in the thermal ensemble give the same expectation value for local observables. The dots in the plot are the values of $M_{ij}/L$, the color intensity being their magnitude. We omit the diagonal matrix elements. The smallest off-diagonal matrix elements are $\approx 10^{-25}\sim e^{-L}$.
It is interesting to investigate the dependence of $M_{ij}$ on the energy difference $E_{i}-E_{j}$. This is illustrated in Fig. 2 (a) plotting $M_{ij}/L$ versus $(E_{i}-E_{j})/\sqrt{L}$, where the $1/\sqrt{L}$ accounts for the fact that in the thermodynamic limit the energy peaks around its thermal value with $\sqrt{L}$ fluctuations. The different colors now correspond to different chain sizes $L\leq 240$, with fixed particle density $M/L=1/4$. The data for different $L$ overlap in the large $L$ limit, suggesting that $M_{ij}\sim L$. The largest off-diagonal matrix elements $M_{ij}$ occur for $E_{j}=E_{i}$, and $M_{ij}$ increases if the eigenstates are farther apart in energy. In Fig. 2 (b) we report with the color scale the number of matrix elements. Clearly, the most abundant values of $M_{ij}$ are at $E_{i}-E_{j}\approx 0$ and peak around $M_{ij}/L\approx-1.5$.

### v2

@@ I Introduction
[…]
@@
where $\overline{E}=\frac{1}{2}(E_{i}+E_{j})$ is the average energy of the eigenstates, $\omega=\left|E_{i}-E_{j}\right|$, $O(\overline{E})$ is a smooth function of the energy $\overline{E}$. In (1), $S(\overline{E})$ is the thermodynamic entropy at $\bar{E}$, $f_{\mathcal{O}}(\omega,\overline{E})$ is a smooth function of its arguments, and $R_{ij}$ is a random variable with mean $\overline{R_{ij}}=0$ and variance $\overline{R_{ij}^{2}}=1$. Eq. (1) implies that in the thermodynamic limit, since the entropy $S$ is extensive in the system size $L$, diagonal matrix elements converge to $O(\overline{E})$, i.e., they become smooth functions of the energy only. Moreover, the variance of diagonal matrix elements decays exponentially with $L$. Analogously, off-diagonal matrix elements are exponentially suppressed with increasing $L$. An intriguing recent result is that the statistics of the off-diagonal matrix elements can be understood in terms of Free Probability Theory [5, 6]. Crucially, Eq. (1) implies local thermalization after quantum quenches in generic, i.e., nonintegrable systems, at least for typical initial states [4]. The ETH scaling (1) has been checked numerically for both diagonal and off-diagonal matrix elements in nonintegrable systems by employing exact diagonalization methods [7, 8, 9, 10, 11, 12, 13, 14]. In integrable models, the presence of an extensive number of conserved charges leads to violations of (1), which have been investigated numerically [15, 16, 17, 18, 19, 20, 21, 22]. Refs. [23, 24] investigated the ETH scaling of diagonal matrix element in interacting integrable systems by exploiting Algebraic Bethe Ansatz (ABA) results [23, 25] for correlation functions, which allow to reach much larger system sizes compared with exact diagonalization. Precisely, Ref. [25] showed that the variance of diagonal matrix elements decays as $L^{-1}$, in contrast with (1).
@@
Very recently, Ref. [26] investigated off-diagonal matrix elements of both local and non-local operators in the Lieb-Liniger model, which describes a one-dimensional system of interacting bosons in the continuum. Ref. [26] showed that the finite-size scaling of off-diagonal matrix elements depends crucially on whether $|E_{i}\rangle$ and $|E_{j}\rangle$ (cf. (1)) are extracted from the same thermodynamic macrostate, i.e., statistical ensemble. Precisely, if $|E_{i}\rangle,|E_{j}\rangle$ are in the same macrostate, off-diagonal matrix elements decay exponentially as [26]
[…]
@@
Quite generically, the PDFs of the logarithm $M_{ij}$ of the absolute value of the matrix elements are well described by Gumbel distributions, in contrast with Ref. [26]. Similar to Ref. [26], however, we observe that the fact that the distribution of $\widetilde{M}_{ij}$ is a Gumbel with $L$-independent parameters is incompatible with the statistics of off-diagonal matrix elements $\langle E_{i}|\mathcal{O}|E_{j}\rangle$ being Gaussian (see Sec.˜III.1.3 for numerical evidence). Next, we consider two-spin operators. We focus on thermal macrostates constructed from eigenstates that do not contain strings. In contrast with one-spin operators, for which we can reach $L\sim 500$, for two-spin operators we only provide numerical data for $L\sim 60$. Despite that, our numerical data confirm the scenario observed for one-spin operators.
@@ II Heisenberg spin chain
The $XXX$ chain Eq.˜5 is a paradigmatic example of Bethe Ansatz solvable many-body system [37, 36, 38, 39]. In the Bethe ansatz treatment of the $XXX$ chain the ferromagnetic state with all the spin up is a “vacuum” state, whereas overturned spins are treated as quasiparticle excitations. Each quasiparticle is labeled by a complex number $\lambda_{i}$, which is called rapidity. Any eigenstate in the sector with fixed number $M$ of quasiparticles, i.e., fixed magnetization, is identified by a set of $M$ rapidities $\lambda_{i}$. The rapidities satisfy nontrivial quantization conditions, encoded in a set of nonlinear coupled algebraic equations as
@@
The Bethe equations (6) admit $2^{L}$ sets of solutions containing both real and complex rapidities. Complex rapidities correspond to bound states of quasiparticles. Although it is in general challenging to obtain the complex solutions of (6), it was already recognized by Bethe [37] that in the thermodynamic limit $L\to\infty$ some crucial simplifications occur. Precisely, the vast majority of complex rapidities form “simple” structures in the complex plane known as Bethe strings. Solutions of (6) in the same string have the same real part (string center) and imaginary parts that differ by $\eta=\mathrm{i}$. A string of length $n$ corresponds to a bound states of $n$ elementary excitations. In generic integrable systems, strings of arbitrary length are allowed.
Within the framework of the string hypothesis, let us denote the rapidities as $\lambda^{j}_{n,\gamma}$, with $\gamma$ labeling the different strings of size $n$, i.e., that have different string centers, and $j$ labeling the rapidities in the same string. For large $L$ one has
@@
where $\delta_{n,\gamma}^{j}$ are the so-called string deviations, which account for the fact that the string hypothesis holds only in the limit $L\to\infty$. In (8) $\lambda_{n,\gamma}$ is the string center, and it is real. Real solutions of the Bethe equations correspond to $n=1$. For the vast majority of the eigenstates of the $XXX$ chain string deviations are suppressed exponentially as $\delta_{j,\gamma}^{j}=\mathcal{O}(e^{-\alpha L})$ (see Ref. [40] for a detailed study of string deviations in the $XXX$ chain). In particular, string deviations are not expected to affect the thermodynamic behavior of the model. By employing (8) in Eq.˜6, after taking the limit $\delta_{n,\gamma}^{j}\to 0$ and considering the logarithm of both members of the Bethe equations, one can write a set of equations for the string centers $\lambda_{n,\gamma}$. One obtains the so-called Bethe-Gaudin-Takahashi equations [36] as
@@
$\Theta_{n,m}=\theta_{|n-m|}+\!\!\!\!\!\!\sum_{r=1}^{(n+m-|n-m|-1)/2}\!\!\!\!\!\!\!\!\!\!\!\!\!2\theta_{|n-m|+2r}+\theta_{n+m},$ (10)
@@
$\Theta_{n,m}=\sum_{r=1}^{n-1}2\theta_{2r}+\theta_{2n},$ (11)
if $n=m$. In Eqs.˜9, 10 and 11 we defined $\theta_{n}(x)=2\arctan(x/n)$. Let us denote by $s_{n}$ the number of strings of length $n$ in the eigenstate. Thus, one has the total number of particles as $M=\sum_{j=1}^{M}js_{j}$. In (9), $I_{n,\gamma}$ are the Bethe-Gaudin-Takahashi (BGT) quantum numbers, and are integer or half-integer numbers for $L-s_{n}$ odd or even, respectively. For the $XXX$ chain the BGT quantum numbers satisfy the bound as [36]
@@ II.1 Macrostates and observables
In the following section, in particular, we will study the statistics of the off-diagonal matrix elements of $E_{i}^{(11)}$ and $E_{i}^{(11)}E_{i+1}^{(22)}$, taken as prototypical examples of one and two-spin operators, respectively. The operators (16) are particularly convenient to compute in the $XXX$ model (5), by virtue of the exact results of Refs. [30, 31]. Indeed, a major breakthrough [30, 31] is that it is possible to rewrite the operators (16) in terms of the matrix elements of the so-called transfer matrix of the $XXX$ chain [38], which is at the heart of the ABA solution of the model. Then, by using the known action of the transfer matrix on the eigenstates of (5), it is possible to obtain the matrix elements. We review the main results of Refs. [30, 31] in Appendix A, reporting the explicit expression for the matrix elements of the operators in (16). Unlike exact diagonalization approaches, the computational cost to evaluate matrix elements via the ABA results reported in appendix˜A scales exponentially only in the size $\ell$ of the support of the operator (here we restrict ourselves to $\ell=1,2$), but polynomially as $O(M^{\ell})$ in the number of particles $M$. This allows us to reach chain sizes $L\sim 500$ which are prohibitively large for exact diagonalization. Still, we should mention that the numerical implementation of the ABA formulas for the matrix elements is nontrivial, because they are derived for an inhomogeneous version of the $XXX$ chain. The homogeneous limit that is required to recover (5) introduces fictitious singularities that have to be removed. Furthermore, the presence of string solutions introduces additional singularities, which have to be dealt with by carefully taking the limit of vanishing string deviations. For the operators with support on more than one site, treating the homogeneous limit and the limit of vanishing string deviations at the same time is very challenging. For this reason, as in Ref. [24], we only consider off-diagonal matrix elements of $E_{i}^{(11)}E_{i+1}^{(22)}$ between eigenstates without bound states. For one-spin operators, since the homogeneous limit is trivial, the limit of vanishing string deviations can be performed analytically. Thus, for one-spin operators we are able to investigate the effect of bound states on the finite-size scaling of matrix elements.
@@ III Numerical results
Figure 1: Logarithm of the matrix elements for the one-spin operator $E^{(11)}_{j}$ with L=56,M=14L=56,\,M=14 as a function of the energy densities $E_{i}/L$ of the two eigenstates (diagonal elements not reported).
Figure 2: Scaling of the logarithm $M_{ij}$ (cf. Eq.˜17) of the matrix elements of the one-spin operator $E_{i}^{(11)}$ as a function of $E_{i}-E_{j}$, with $E_{i},E_{j}$ the energy of the two eigenstates. (a) Scatter plot of the logarithm $M_{ij}$ rescaled by $L$ as a function of $(E_{i}-E_{j})\sqrt{L}$ for $L=56,80,120,160,$ and $240$ and fixed $M=L/4$. (b) Two-dimensional histogram of the rescaled logarithm $M_{ij}/L$ as a function of $(E_{i}-E_{j})/\sqrt{L}$.
@@
[…]
In Fig.˜1 we overview the behavior of $M_{ij}$ for the one-spin operator $E^{(11)}_{i}$ (cf. (16)) in the $XXX$ chain with $L=56$, $M=L/4$. The data shown in the figure are for the infinite-temperature macrostate. Precisely, by employing the method of Ref. [42] we generate an ensemble of $\sim 10^{3}$ eigenstates of the $XXX$ chain distributed with the infinite-temperature Gibbs probability. As discussed earlier, we fix the magnetization $M/L=1/4$ and consider only the eigenstates that correspond to solutions of the Bethe equations without infinite rapidities. On the two axes we show the energy density $E_{1}/L$ and $E_{2}/L$ of the eigenstates. Here we have $E_{1}/L,E_{2}/L\in[-0.3,-0.1]$. Notice that in the thermodynamic limit $L,M\to\infty$, the fluctuations of the energy density are suppressed as $1/\sqrt{L}$, i.e., the energy density peaks at the thermal value at the density $M/L=1/4$. Again, this reflects that in the thermodynamic limit all the eigenstates in the thermal ensemble give the same expectation value for local observables. The dots in the plot are the values of $M_{ij}/L$, the color intensity being their magnitude. We omit the diagonal matrix elements. The smallest off-diagonal matrix elements are $\approx 10^{-25}\sim e^{-L}$.
It is interesting to investigate the dependence of $M_{ij}$ on the energy difference $E_{i}-E_{j}$. This is illustrated in Fig.˜2 (a) plotting $M_{ij}/L$ versus $(E_{i}-E_{j})/\sqrt{L}$, where the $1/\sqrt{L}$ accounts for the fact that in the thermodynamic limit the energy peaks around its thermal value with $\sqrt{L}$ fluctuations. The different colors now correspond to different chain sizes $L\leq 240$, with fixed particle density $M/L=1/4$. The data for different $L$ overlap in the large $L$ limit, suggesting that $M_{ij}\sim L$. The largest off-diagonal matrix elements $M_{ij}$ occur for $E_{j}=E_{i}$, and $M_{ij}$ increases if the eigenstates are farther apart in energy. In Fig.˜2 (b) we report with the color scale the number of matrix elements. Clearly, the most abundant values of $M_{ij}$ are at $E_{i}-E_{j}\approx 0$ and peak around $M_{ij}/L\approx-1.5$.